A 20 μF capacitor is joined to a battery of emf 6.0 V through a resistance of 100 Ω. Find the charge on the capacitor 2.0 ms after the connections are made.

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. C = 20 * 10^–6 F, E = 6 V, R = 100 Ω t = 2 * 10^–3 sec q = EC (1 – e^–t/RC) = 6 * 20 * 10^-6 (1 – e^-2*10^-3/100*20*10^-6) = 12 x 10^–5 (1 – e^–1) = 7.12 * 0.63 * 10^–5 = 7.56 * 10^–5 = 75.6 * 10^–6 = 76 μc.